OpenStudy (anonymous):

Find all the zeroes of the equation. -3x^4+27x^2+1200=0

OpenStudy (mertsj):

Factor out -3 first

OpenStudy (anonymous):

ok so -3 (x^4+9x^2+400)

OpenStudy (mertsj):

-3(x^4-9x^2-400)

OpenStudy (mertsj):

Now find two factors of -400 that add up to -9

OpenStudy (mertsj):

I'll give you a hint: One of them is -25

OpenStudy (anonymous):

16,-25

OpenStudy (mertsj):

Now write the factors of x^4-9x^2-400

OpenStudy (anonymous):

im confused

OpenStudy (mertsj):

Fill in the blanks: x^4-9x^2-25=(x^2- )(x^2+ )

OpenStudy (anonymous):

(x^2-16)(x^2-25)

OpenStudy (mertsj):

So now we have this: -3(x^2-25)(x^2+16)=0

OpenStudy (anonymous):

ok

OpenStudy (mertsj):

Set each factor equal to 0 and solve each equation.

OpenStudy (anonymous):

-5,5?

OpenStudy (mertsj):

Those are two of the answers. Now do the other equation.

OpenStudy (anonymous):

-4,4

OpenStudy (mertsj):

\[x^2+16=0\] \[x^2=-16\] \[\sqrt{x^2}=\pm \sqrt{-16}\]

OpenStudy (mertsj):

\[x=\pm \sqrt{(-1)(16)}\] \[x=\pm \sqrt{-1}\sqrt{16}\] \[x=\pm i(4)= \pm 4 i\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

is that the final answer

OpenStudy (mertsj):

Those are the final 4 answers. Since your equation has x^4, there should be 4 answers.

OpenStudy (anonymous):

ok thank you!

OpenStudy (mertsj):

yw